Find the frequency of the electron in an orbit of radius r if velocity is v.
 

Grade:11

1 Answers

Arun
4 years ago
Angular velocity, w= V/r

Relation between angular velocity and frequency is w = 2×pi×frequency

Therefore f= V/(2×pi×r)